# DiG-Net: Enhancing Quality of Life through Hyper-Range Dynamic Gesture Recognition in Assistive Robotics

### v1

@@ Abstract
Dynamic hand gestures play a pivotal role in assistive human-robot interaction (HRI), facilitating intuitive, non-verbal communication, particularly for individuals with mobility constraints or those operating robots remotely. Current gesture recognition methods are mostly limited to short-range interactions, reducing their utility in scenarios demanding robust assistive communication from afar. In this paper, we introduce a novel approach designed specifically for assistive robotics, enabling dynamic gesture recognition at extended distances of up to 30 meters, thereby significantly improving accessibility and quality of life. Our proposed Distance-aware Gesture Network (DiG-Net) effectively combines Depth-Conditioned Deformable Alignment (DADA) blocks with Spatio-Temporal Graph modules, enabling robust processing and classification of gesture sequences captured under challenging conditions, including significant physical attenuation, reduced resolution, and dynamic gesture variations commonly experienced in real-world assistive environments. We further introduce the Radiometric Spatio‐Temporal Depth Attenuation Loss (RSTDAL), shown to enhance learning and strengthen model robustness across varying distances. Our model demonstrates significant performance improvement over state-of-the-art gesture recognition frameworks, achieving a recognition accuracy of 97.3% on a diverse dataset with challenging hyper-range gestures. By effectively interpreting gestures from considerable distances, DiG-Net significantly enhances the usability of assistive robots in home healthcare, industrial safety, and remote assistance scenarios, enabling seamless and intuitive interactions for users regardless of physical limitations.
keywords:
@@
Dynamic Gesture Recognition , Assistive Robotics , Human-Robot Interaction , Hyper-Range Perception , Graph Transformer
journal: Computer Vision and Image Understanding
@@
[mit] organization=Department of Mechanical Engineering, Massachusetts Institute of Technology (MIT), addressline=77 Massachusetts Ave, city=Cambridge, postcode=02139, state=MA, country=USA
[…]
[tau] organization=School of Mechanical Engineering, Tel-Aviv University, addressline=P.O. Box 39040, city=Tel-Aviv, postcode=6997801, state=, country=Israel
1 Introduction
The growing number of individuals living with disabilities and requiring assistance has created a pressing demand for assistive technologies that enhance users’ independence, safety, and quality of life [1]. Among these, assistive robotic systems are increasingly integrated into environments where intuitive, nonverbal communication is essential for enabling natural interaction with individuals of varied abilities. Gesture-based interaction is particularly important in scenarios where speech is not an option. To contextualize our contribution, Table 1 presents a comparative overview of recent systems in this area, outlining their target users, sensing modalities, application domains, and level of human involvement. Our method is the only one to support dynamic gesture recognition at hyper-range distances, defined here as up to 30 meters, and to operate reliably in both indoor and outdoor environments, making it uniquely suited for real-world assistive deployment. Recent research has catalyzed a new generation of assistive systems capable of perceiving complex environments and interacting with humans in context-aware, natural ways [2, 3, 4]. Equally important, human-in-the-loop approaches are being developed to ensure these systems operate safely and reliably alongside people. These methods incorporate human feedback, predictive planning, and adaptive control to support trust and safety during close interactions [5, 6, 7]. To enable practical deployment, recent studies stress the importance of user-centered design and usability evaluation, ensuring assistive robots are intuitive, reliable, and effective [8, 9]. A key area of research supporting these goals is human-robot interaction (HRI), which focuses on enabling smooth, natural collaboration between people and robots, both in assistive contexts and everyday life [10, 11]. This reflects a broader shift from viewing robots as tools to recognizing them as partners that augment and support human capabilities [12].
Table 1: Comparative Summary of Gesture-Based Assistive Robotics Systems
Paper Population Env Modality Application Range Human-in-the-Loop Boboc et al. [13] General users Indoor RGB-D Home robot control Short Yes Haseeb et al. [14] Motor-impaired Indoor IMU Head-based robot control N/A Yes Rudigkeit et al. [15] Tetraplegic Indoor IMU Robot arm use N/A Yes Yang et al. [16] Upper-limb impaired Indoor EMG Mobile manipulation Long Yes Ababneh et al. [17] Elderly, wheelchair Indoor RGB-D Arm control via gesture Short Yes Neto et al. [18] Factory workers Indoor IMU Assembly support Mid Partial Werner et al. [19] Elderly Indoor RGB-D Bathing assistant Short Partial Muñoz et al. [20] Older adults Indoor RGB-D Rehab & exercises Short Partial Oudah et al. [21] Speech-impaired elderly Indoor RGB-D Emergency alerting Mid Yes Bandara et al. [22] Wheelchair users Indoor RGB-D Wheelchair navigation Short Yes This work General users Both RGB Remote HRI guidance Hyper Yes
Intuitive interaction mechanisms are essential to enable non-expert users to effectively communicate their intentions to robots. Hand and arm gestures are intuitive forms of nonverbal communication, widely used in daily human interactions, making them well-suited for HRI [23, 24, 25]. Gesture-based interactions reduce the need for complex verbal or physical interfaces, improve user experience, and allow for intuitive and rapid communication with robots, even for users with no technical expertise [26]. With gestures, a user can convey nonverbal and simple commands even from a long distance without the need to shout. For instance, a user may direct robot movements with pointing gestures [27].
Figure 1: Demonstration of a user instructing a robot to go back by sweeping an open palm forward and backward, from a hyper-range distance. In addition to the low-resolution view of the user’s hand, the robot may confuse the dynamic gesture with the static stop gesture.
Research in the field of gesture recognition has made significant progress, particularly in recognizing static gestures [28, 29, 30]. However, most gesture recognition approaches are limited to short-range interactions, typically within a few meters [31, 26, 24]. Moreover, gesture recognition methods, often described to be effective in the long range, are typically limited to around seven meters [32, 33]. However, this limitation hinders their applicability in real-world scenarios that demand longer-range HRI. Recognizing gestures from a truly long distance can significantly expand the potential applications of robots in environments such as public spaces, industrial settings, and emergencies, where natural, non-contact interactions are necessary [34, 35]. However, one of the key challenges in achieving effective long-range gesture recognition is the degradation of visual information due to factors such as reduced resolution, lighting variations, and occlusions [36, 37].
While our recent work [38] successfully demonstrated ultra-range recognition of static hand gestures using a simple web camera at distances up to 25 meters, recognizing dynamic gestures presents additional, significant challenges. Static gesture recognition relies primarily on spatial features captured in a single frame, whereas dynamic gesture recognition demands the extraction and analysis of temporal sequences across multiple frames, significantly increasing complexity due to motion blur, occlusions, and ambiguity between similar gestures. The transition from static to dynamic gestures represents a fundamental shift, requiring entirely new approaches rather than incremental extensions of existing methods. Dynamic gesture recognition at hyper-range presents a unique challenge that demands the fusion of spatial and temporal information across multiple scales and distances. At such long ranges, the visual signal of the performer is severely diminished, the region of interest occupies only a small portion of the frame, resulting in low-resolution observations, and environmental noise (e.g., background clutter, illumination changes, atmospheric effects) can significantly degrade each frame. In this scenario, temporal cues become as crucial as spatial cues for reliable recognition. Fast, subtle hand motions (e.g., a quick finger wave) might be nearly imperceptible in any single low-resolution frame, yet can be captured by analyzing differences across consecutive frames. Conversely, the overall spatial configuration of a gesture (e.g., an arm extended forward) may be visible but ambiguous without temporal context to distinguish its meaning. This interplay between space and time at extreme distances requires an architecture that can simultaneously capture fine, rapid movements and aggregate information over longer temporal intervals.
Figure 2: Overview of the proposed DiG-Net framework for Hyper-range dynamic hand gesture recognition. The framework starts with feature extraction using ResNet, followed by frame reduction using K-Means clustering. User detection is performed using YOLOv3, with the output frames resized to $224\times 224$. The reduced frames are processed through the DiG-Net to robustly recognize gestures from RGB videos at distances up to 30 meters. The model integrates Depth-Conditioned Deformable Alignment (DADA), Spatio-Temporal Graph (STG) modules, and Graph Transformer encoders. Finally, a classification head is employed to acquire the gesture class.
Dynamic gestures provide valuable context for understanding actions and systems, especially when explaining processes that unfold over time [39]. They are particularly effective for conveying spatial context to a task command [40]. For instance, a gesture instructing a robot to move backward would usually include an open palm swept back and forth toward it, as demonstrated in Figure 1. However, a snapshot of the gesture may be recognized as a static stop instruction if not processed correctly. Several studies have attempted to address dynamic gesture recognition by utilizing various data modalities. RGB-Depth (RGB-D) cameras are often used to recognize dynamic gestures but are limited to indoor environments [41, 42, 43, 44, 45]. Approaches using simple RGB cameras were also proposed while shown to function only indoors [32, 46, 47]. Table 2 summarizes the prominent work on dynamic gesture recognition with visual perception, highlighting the limitations of current approaches, which are typically restricted to indoor environments and short-range interactions (within 7 meters). In it also worth noting on-body sensors in wearable devices that instantly recognize hand gestures [48, 49]. While they can offer high recognition rates, wearable devices often require specialized and expensive hardware, limiting their accessibility to occasional users. Additionally, biometric approaches may not generalize well to new users, requiring additional data collection for each individual. On the other hand, visual perception allows for the observation of any user.
Table 2: State-of-the-art comparison of dynamic gesture recognition methods using visual perception
Paper Camera Range Indoor Outdoor
Zhou et al. [32] RGB $\leq 7m$
Xu et al. [41] RGB-D $<1m$
Ma et al. [42] RGB-D $\leq 2m$
Kabir et al. [43] RGB-D $<1m$
Bokstaller et al. [44] RGB-D -
Gao et al. [45] RGB-D $<2m$
Yi et al. [46] RGB $\leq 6m$
dos Santos et al. [47] RGB -
Tran et al. [50] 5$\times$×RGB-D $<2m$
Wu et al. [51] RGB-D $<2m$
Proposed method RGB $\leq 30m$
@@ 1 Introduction
While the above methods show feasibility in controlled indoor settings, they often lack scalability and robustness for outdoor environments or hyper-range interactions. Their reliance on specialized hardware, such as a depth cameras, can increase cost and complexity, limiting their practical applicability [52]. Recent approaches utilizing multimodal data, including RGB, depth and human-pose (i.e., skeleton) information, have demonstrated potential for dynamic gesture recognition [50, 51]. However, these methods are often also constrained by limited operational range and the need for specialized equipment. This makes them less suitable for scenarios requiring long-range interactions or deployment in large, open spaces. To the best of the authors’ knowledge, no work has addressed the problem of recognizing dynamic gestures at distances farther than seven meters.
@@
In addition to capturing temporal features of dynamic gestures, the challenge is further compounded by the difficulty of perceiving spatial motions at far distances. In this paper, we address the problem of dynamic gesture recognition at hyper-range distance by only using a simple RGB camera. Specifically, we propose the Distance-aware Gesture Network (DiG-Net), illustrated in Figure 2, which comprises a novel Depth-Conditioned Deformable Alignment (DADA) block, seamlessly fused with Spatio-Temporal Graph (STG) modules [53]. Unlike existing approaches, DiG-Net adaptively warps feature maps based on per-pixel depth estimates to compensate for physical attenuation and defocus blur, while graph-based temporal modeling captures dynamic gesture patterns across frames. This unified design ensures robust feature alignment and temporal coherence, enabling accurate classification of subtle gestures at distances up to 30 meters under low-resolution and noisy conditions [54], which are common in hyper-range gesture recognition scenarios [37, 55].
@@
To encourage the model to perform better at hyper-range distances, we introduce the Radiometric Spatio‐Temporal Depth Attenuation Loss (RSTDAL). This specialized loss function incorporates both Beer–Lambert attenuation and defocus weighting to adaptively adjust each training sample’s contribution based on its distance from the camera, thereby strengthening recognition accuracy for gestures performed at greater distances. The proposed DiG-Net trained using the RSTDAL loss yields a robust and generalizable recognition for dynamic gestures in both indoor and outdoor environments. The model addresses key challenges in hyper-range gesture recognition, such as the degradation of visual information due to reduced resolution and lighting variations [56]. Furthermore, we introduce two novel metrics specifically designed to evaluate gesture recognition performance at long distances, including the stability of the recognition over time. Our approach is evaluated using these metrics, along with standard evaluation metrics.
@@
By enabling accurate dynamic gesture recognition at hyper-range distances, we significantly expand the potential applications of robots in environments such as public spaces, industrial settings, and emergency situations, where natural, non-contact interactions are necessary. Furthermore, our approach may enhance the ability of robots to interact with multiple users simultaneously, improving their utility in crowded or dynamic environments [57, 58]. Our method can also be applied in various fields, including search and rescue, drone operations and service robotics. Additionally, its applications may extend to space exploration, law enforcement and entertainment.
2 Methods
2.1 Problem Formulation

### v2

# DiG-Net: Enhancing Human–Robot Interaction through Hyper-Range Dynamic Gesture Recognition in Assistive Robotics

@@ Abstract.
Dynamic hand gestures play a pivotal role in assistive human-robot interaction (HRI), facilitating intuitive, non-verbal communication, particularly for individuals with mobility constraints or those operating robots remotely. Current gesture recognition methods are mostly limited to short-range interactions, reducing their utility in scenarios demanding robust assistive communication from afar. In this paper, we present DiG-Net, the first dynamic gesture recognition framework enabling robust operation at hyper-range distances of up to 30 meters, specifically designed for assistive robotics to enhance accessibility and improve quality of life. Our proposed Distance-aware Gesture Network (DiG-Net) effectively combines Depth-Conditioned Deformable Alignment (DADA) blocks with Spatio-Temporal Graph modules, enabling robust processing and classification of gesture sequences captured under challenging conditions, including significant physical attenuation, reduced resolution, and dynamic gesture variations commonly experienced in real-world assistive environments. We further introduce the Radiometric Spatio‐Temporal Depth Attenuation Loss (RSTDAL), shown to enhance learning and strengthen model robustness across varying distances. Our model demonstrates significant performance improvement over state-of-the-art gesture recognition frameworks, achieving a recognition accuracy of 97.3% on a diverse dataset with challenging hyper-range gestures. By effectively interpreting gestures from considerable distances, DiG-Net significantly enhances the usability of assistive robots in home healthcare, industrial safety, and remote assistance scenarios, enabling seamless and intuitive interactions for users regardless of physical limitations.
@@
Human-Robot Interaction, Assistive Robotics, Dynamic Gesture Recognition, Hyper-Range Perception, Human-Centered Design, Graph Transformer
copyright: acmlicensedjournal: THRIjournalyear: 2025ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Human robot interaction (HRI)$c$s: Computing methodologies Machine learning approachesccs: Computing methodologies Computer vision problemsccs: Human-centered computing Accessibility systems and tools
1. Introduction
The growing number of individuals living with disabilities and requiring assistance has created a pressing demand for assistive technologies that enhance users’ independence, safety, and quality of life (Nanavati et al., 2023). Among these, assistive robotic systems are increasingly integrated into environments where intuitive, nonverbal communication is essential for enabling natural interaction with individuals of varied abilities. Gesture-based interaction is particularly important in scenarios where speech is not an option. To contextualize our contribution, Table 1 presents a comparative overview of recent systems in this area, outlining their target users, sensing modalities, application domains, and level of human involvement. Recent developments in gesture-based assistive systems have improved short-range communication, yet reliable operation at long distances remains largely unexplored. This work addresses that gap by enabling robust dynamic gesture recognition suitable for real-world assistive deployment. Recent research has catalyzed a new generation of assistive systems capable of perceiving complex environments and interacting with humans in context-aware, natural ways (Leal-Taixé and Roth, 2019; Robinson et al., 2023; Drolshagen et al., 2023). Equally important, human-in-the-loop approaches are being developed to ensure these systems operate safely and reliably alongside people. These methods incorporate human feedback, predictive planning, and adaptive control to support trust and safety during close interactions (Lasota et al., 2017; Safeea and Neto, 2019; Losey et al., 2018). To enable practical deployment, recent studies stress the importance of user-centered design and usability evaluation, ensuring assistive robots are intuitive, reliable, and effective (Frennert et al., 2024; Shourmasti et al., 2021). A key area of research supporting these goals is HRI, which focuses on enabling smooth, natural collaboration between people and robots, both in assistive contexts and everyday life (Goodrich et al., 2008; Maurtua et al., 2017). This reflects a broader shift from viewing robots as tools to recognizing them as partners that augment and support human capabilities (Matarić, 2017).
Beyond these technological advances, recent literature in HRI emphasizes that social perception, trust, and intuitiveness are equally essential for effective collaboration. Nonverbal cues, such as gestures, posture, and motion timing, play a fundamental role in how humans interpret robot behavior and establish mutual understanding. Urakami and Seaborn (Urakami and Seaborn, 2023) highlighted that natural gesture expressiveness and responsiveness significantly affect user comfort and perceived agency during interaction. Similarly, Robinson et al. (Robinson et al., 2023) stressed the growing importance of robotic vision that supports collaboration through real-time, context-aware perception. As Mataric (Matarić, 2017) argued in her influential review on socially assistive robotics, the ultimate goal is to move from automation toward human augmentation, developing robots that enhance human capability rather than replace it. Framing DiG-Net within this paradigm positions it not merely as a gesture recognition model, but as a medium for intuitive, trustworthy, and socially aware communication between humans and assistive robots.
Table 1. Comparative summary of gesture-based assistive robotics systems.
Paper Population Env. Modality Application Range Human-in-the-loop Boboc et al. (Boboc et al., 2015) General users Indoor RGB-D Home robot control Short Yes Haseeb et al. (Haseeb et al., 2018) Motor-impaired Indoor IMU Head-based robot control N/A Yes Rudigkeit et al. (Rudigkeit and Gebhard, 2019) Tetraplegic Indoor IMU Robot arm use N/A Yes Yang et al. (Yang et al., 2025) Upper-limb impaired Indoor EMG Mobile manipulation Long Yes Ababneh et al. (Ababneh et al., 2018) Elderly, wheelchair Indoor RGB-D Arm control via gesture Short Yes Neto et al. (Neto et al., 2019) Factory workers Indoor IMU Assembly support Mid Partial Werner et al. (Werner et al., 2020) Elderly Indoor RGB-D Bathing assistant Short Partial Muñoz et al. (Muñoz et al., 2021) Older adults Indoor RGB-D Rehab & exercises Short Partial Oudah et al. (Oudah et al., 2020) Speech-impaired elderly Indoor RGB-D Emergency alerting Mid Yes Bandara et al. (Bandara et al., 2020) Wheelchair users Indoor RGB-D Wheelchair navigation Short Yes This work General users Both RGB Remote HRI guidance Hyper Yes
Intuitive interaction mechanisms are essential to enable non-expert users to effectively communicate their intentions to robots. Hand and arm gestures are intuitive forms of nonverbal communication, widely used in daily human interactions, making them well-suited for HRI (Gao et al., 2020; Wang et al., 2022; Urakami and Seaborn, 2023). Gesture-based interactions reduce the need for complex verbal or physical interfaces, improve user experience, and allow for intuitive and rapid communication with robots, even for users with no technical expertise (Wachs et al., 2011). With gestures, a user can convey nonverbal and simple commands even from a long distance without the need to shout. For instance, a user may direct robot movements with pointing gestures (Bamani et al., 2025).
Figure 1. Demonstration of a user instructing a robot to go back by sweeping an open palm forward and backward, from a hyper-range distance. In addition to the low-resolution view of the user’s hand, the robot may confuse the dynamic gesture with the static stop gesture.
A person standing far from a mobile robot performs a dynamic back-and-forth palm gesture. The image shows how low resolution at a long distance can cause the robot to misinterpret the motion as a static stop signal.
Research in the field of gesture recognition has made significant progress, particularly in recognizing static gestures (Zhang et al., 2020; Yu et al., 2021, 2022). However, most gesture recognition approaches are limited to short-range interactions, typically within a few meters (Nickel and Stiefelhagen, 2007; Wachs et al., 2011; Wang et al., 2022). Moreover, gesture recognition methods, often described to be effective in the long range, are typically limited to around seven meters (Zhou et al., 2021; Liang et al., 2024). However, this limitation hinders their applicability in real-world scenarios that demand longer-range HRI. Recognizing gestures from a truly long distance can significantly expand the potential applications of robots in environments such as public spaces, industrial settings, and emergencies, where natural, non-contact interactions are necessary (Kim and Moon, 2014; Canal et al., 2015). However, one of the key challenges in achieving effective long-range gesture recognition is the degradation of visual information due to factors such as reduced resolution, lighting variations, and occlusions (Shen et al., 2019; Pla et al., 2018).
While our recent work (Bamani et al., In press, 2024) successfully demonstrated ultra-range recognition of static hand gestures using a simple web camera at distances up to 25 meters, recognizing dynamic gestures presents additional, significant challenges. Static gesture recognition relies primarily on spatial features captured in a single frame, whereas dynamic gesture recognition demands the extraction and analysis of temporal sequences across multiple frames. Prior works have addressed challenges such as occlusion and dynamic ambiguity in short-range settings, for example, by using shape-based descriptors or customized tracking to handle missing or discontinuous cues (Xing et al., 2019; Han et al., 2018; Jain et al., 2023). However, these approaches typically assume that the hand region is sufficiently resolved and that reliable hand-configuration cues are available, assumptions that can break down at long distances where the gesture occupies only a small portion of the image. As a result, methods designed for short-range sensing may not transfer directly to long-distance, RGB-only deployment without explicitly accounting for compounded degradations. Dynamic gesture recognition at hyper-range presents a unique challenge that demands the fusion of spatial and temporal information across multiple scales and distances. At such long ranges, the visual signal of the performer is severely diminished, the region of interest occupies only a small portion of the frame, resulting in low-resolution observations, and environmental noise (e.g., background clutter, illumination changes, atmospheric effects) can significantly degrade each frame. In this scenario, temporal cues become as crucial as spatial cues for reliable recognition. Fast, subtle hand motions (e.g., a quick finger wave) might be nearly imperceptible in any single low-resolution frame, yet can be captured by analyzing differences across consecutive frames. Conversely, the overall spatial configuration of a gesture (e.g., an arm extended forward) may be visible but ambiguous without a temporal context to distinguish its meaning. This interplay between space and time at extreme distances requires an architecture that can simultaneously capture fine, rapid movements and aggregate information over longer temporal intervals.
Figure 2. Overview of the proposed DiG-Net framework for hyper-range dynamic hand gesture recognition. The model combines Depth-Conditioned Deformable Alignment (DADA), Spatio-Temporal Graph (STG) modules, and Graph Transformer encoders to recognize gestures from RGB videos at distances up to 30 meters.
A schematic diagram illustrating the DiG-Net architecture for hyper-range dynamic gesture recognition. The pipeline begins with feature extraction via ResNet, followed by K-means-based frame reduction and YOLOv3 user detection. Processed frames are fed into the DiG-Net, which integrates DADA blocks for depth-conditioned alignment, STG modules for temporal reasoning, and Graph Transformer encoders for spatial-temporal fusion, ending with a classification head outputting gesture classes.
Dynamic gestures provide valuable context for understanding actions and systems, especially when explaining processes that unfold over time (Kang and Tversky, 2016). They are particularly effective for conveying spatial context to a task command (Clough and Duff, 2020). For instance, a gesture instructing a robot to move backward would usually include an open palm swept back and forth toward it, as demonstrated in Figure 1. However, a snapshot of the gesture may be recognized as a static stop instruction if not processed correctly. Several studies have attempted to address dynamic gesture recognition by utilizing various data modalities. RGB-Depth (RGB-D) cameras are often used to recognize dynamic gestures but are limited to indoor environments (Xu et al., 2015; Ma and Peng, 2018; Kabir et al., 2019; Bokstaller and Improta, 2021; Gao et al., 2022). Approaches using simple RGB cameras were also proposed while shown to function only indoors (Zhou et al., 2021; Yi et al., 2018; dos Santos et al., 2020). Table 2 summarizes the prominent work on dynamic gesture recognition with visual perception, highlighting the limitations of current approaches, which are typically restricted to indoor environments and short-range interactions (within 7 meters). In it also worth noting on-body sensors in wearable devices that instantly recognize hand gestures (Pyun et al., 2023; Tchantchane et al., 2023). While they can offer high recognition rates, wearable devices often require specialized and expensive hardware, limiting their accessibility to occasional users. Additionally, biometric approaches may not generalize well to new users, requiring additional data collection for each individual. On the other hand, visual perception allows for the observation of any user.
Table 2. State-of-the-art comparison of dynamic gesture recognition methods using visual perception.
@@ 1. Introduction
While the above methods show feasibility in controlled indoor settings, they often lack scalability and robustness for outdoor environments or hyper-range interactions. Their reliance on specialized hardware, such as depth cameras, can increase cost and complexity, limiting their practical applicability (Ji et al., 2012). Recent approaches utilizing multimodal data, including RGB, depth, and human-pose (i.e., skeleton) information, have demonstrated potential for dynamic gesture recognition (Tran et al., 2019; Wu et al., 2016). However, these methods are often also constrained by limited operational range and the need for specialized equipment. This makes them less suitable for scenarios requiring long-range interactions or deployment in large, open spaces. To the best of the authors’ knowledge, no work has addressed the problem of recognizing dynamic gestures at distances farther than seven meters.
@@
In addition to capturing temporal features of dynamic gestures, the challenge is further compounded by the difficulty of perceiving spatial motions at far distances. In this paper, we address the problem of dynamic gesture recognition at hyper-range distance by only using a simple RGB camera. Specifically, we propose the Distance-aware Gesture Network (DiG-Net), illustrated in Figure 2, which comprises a novel Depth-Conditioned Deformable Alignment (DADA) block, seamlessly fused with Spatio-Temporal Graph (STG) modules (Yan et al., 2018). Unlike existing approaches, DiG-Net adaptively warps feature maps based on per-pixel depth estimates to compensate for physical attenuation and defocus blur, while graph-based temporal modeling captures dynamic gesture patterns across frames. This unified design ensures robust feature alignment and temporal coherence, enabling accurate classification of subtle gestures at distances up to 30 meters under low-resolution and noisy conditions (Godard et al., 2019), which are common in hyper-range gesture recognition scenarios (Pla et al., 2018; Ren et al., 2011).
@@
To encourage the model to perform better at hyper-range distances, we introduce the Radiometric Spatio‐Temporal Depth Attenuation Loss (RSTDAL). This specialized loss function incorporates both Beer–Lambert attenuation and defocus weighting to adaptively adjust each training sample’s contribution based on its distance from the camera, thereby strengthening recognition accuracy for gestures performed at greater distances. The proposed DiG-Net trained using the RSTDAL loss yields a robust and generalizable recognition for dynamic gestures in both indoor and outdoor environments. The model addresses key challenges in hyper-range gesture recognition, such as the degradation of visual information due to reduced resolution and lighting variations (Gao et al., 2024). Furthermore, we introduce two novel metrics specifically designed to evaluate gesture recognition performance at long distances, including the stability of the recognition over time. Our approach is evaluated using these metrics, along with standard evaluation metrics.
@@
By extending reliable gesture recognition to hyper-range distances, this work aims to broaden the accessibility and usability of assistive robotic systems.
2. Methods
2.1. Problem Formulation